OpenStudy (anonymous):

Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20 cospi over 15 + 30, where H represents the height of the person above the ground in feet at t seconds.

OpenStudy (anonymous):

How long does the Ferris wheel take to make one complete revolution?

OpenStudy (anonymous):

@aum

OpenStudy (aum):

In y = Acos(Bx) + C the period = 2 * pi / B How long does the Ferris wheel take to make one complete revolution? They are asking for the period. Identify B in the cosine function and find 2 * pi / B